### v1

# LLMs Love Python: A Study of LLMs’ Bias for Programming Languages and Libraries

Abstract.
Programming language and library choices are crucial to software reliability and security. Poor or inconsistent choices can lead to increased technical debt, security vulnerabilities, and even catastrophic failures in safety-critical systems. As Large Language Models (LLMs) play an increasing role in code generation, it is essential to understand how they make these decisions. However, little is known about their preferences when selecting programming languages and libraries for different coding tasks. To fill this gap, this study provides the first in-depth investigation into LLM preferences for programming languages and libraries used when generating code. We assess the preferences of eight diverse LLMs by prompting them to complete various coding tasks, including widely-studied benchmarks and the more practical task of generating the initial structural code for new projects (a crucial step that often determines a project’s language or library choices).
Our findings reveal that LLMs heavily favour Python when solving language-agnostic problems, using it in 90%-97% of cases for benchmark tasks. Even when generating initial project code where Python is not a suitable language, it remains the most-used language in 58% of instances. Moreover, LLMs contradict their own language recommendations in 83% of project initialisation tasks, raising concerns about their reliability in guiding language selection. Similar biases toward well-established libraries further create serious discoverability challenges for newer open-source projects. These results highlight the need to improve LLMs’ adaptability to diverse programming contexts and to develop mechanisms for mitigating programming language and library bias.
1. Introduction
The last few years have seen Large Language Models (LLMs) make huge advances (naveedComprehensiveOverviewLarge2024a), particularly excelling in code generation tasks (jiangSurveyLargeLanguage2024). Widely accessible tools like ChatGPT (openaiGPT4TechnicalReport2024) and GitHub Copilot (chenEvaluatingLargeLanguage2021; GitHubCopilotYour2024) enable developers of all experience levels to leverage generative AI in their workflow, leading to a significant impact on newly produced code (staffSurveyAIWave2024; lanzStabilityAICEO2023; mckenna25GooglesCode2024).
While LLMs have demonstrated impressive performance, little is known about their preferences when selecting programming languages and libraries for various coding tasks. Existing research has primarily focused on evaluating and improving their correctness, security, and efficiency in generating code (chenSurveyEvaluatingLarge2024), but the decision-making process behind language and library selection remains unexplored. This gap is particularly significant as LLMs are increasingly used in real-world software development workflows, where informed programming language and library choices are vital for ensuring project success. If these LLMs exhibit systematic biases or inconsistent recommendations, they may mislead developers into making flawed decisions, potentially compromising software reliability, security, and maintainability. In critical applications—such as healthcare, finance, and autonomous systems—these biases could lead to severe safety risks, security vulnerabilities, and system failures, endangering both users and infrastructure.
LLMs may develop strong preferences for certain programming languages or libraries due to various factors, such as the distribution of languages in their training data (wangExploringMultiLingualBias2024), the prevalence of certain libraries in open-source repositories, or biases introduced during model training and fine-tuning (zieglerFineTuningLanguageModels2020). For example, if an LLM is trained on codebases where a particular language dominates, it may exhibit a strong inclination toward generating solutions in that language, even when alternative languages would be more appropriate for the task. As LLMs are increasingly integrated into software development workflows, understanding these preferences is essential not only for assessing their reliability and applicability across diverse programming scenarios but also for developing strategies to mitigate potential biases and improve their adaptability.
To fill this gap, this paper provides the first empirical investigation into the programming language and library preferences of LLMs when writing code. As the first paper to consider and discuss such preferences, we focus on two primary aspects of coding: programming languages, as they are intrinsic to coding; and libraries, as they provide essential pre-built functionality, and play a crucial role in modern software development (somervilleSoftwareEngineeringGlobal2016).
We measure LLM preferences by using prompts that require code generation without explicitly specifying the programming language or library — an approach that reflects common real-world scenarios, particularly for low-code users and developers initiating new projects (rasnayakaEmpiricalStudyUsage2024). The responses are then processed, with languages and libraries extracted for further analysis. To ensure a broad understanding of LLM preferences, eight widely-used LLMs are chosen for the study, covering both open- and closed-source and a range of sizes: GPT-4o (openaiGPT4TechnicalReport2024), GPT-3.5 (andrewpengGPT35TurboFinetuning), Claude3.5 (Haiku and Sonnet) (anthropicClaude3Model2024), Llama3.2 (grattafioriLlama3Herd2024), Qwen2.5 (huiQwen25CoderTechnicalReport2024), DeepSeekLLM (deepseek-aiDeepSeekLLMScaling2024) and Mistral7b (jiangMistral7B2023).
For both programming languages and libraries, we examine LLM preferences in two key scenarios: (1) solving problems from widely-studied benchmarks, and (2) generating the initial code for a new project, a crucial phase where foundational technology choices—such as programming language and library selection—are made. Additionally, we examine whether all LLMs share similar preferences. We also assess their internal consistency by checking if the languages and libraries they recommend match those they actually use when generating code for the same task.
Our experiments lead to the following primary observations: 1) All LLMs we study exhibit a significant bias toward Python as their preferred programming language. To our surprise, for scenarios where Python is not a suitable language, Python remains the dominant choice as the most-used language in 58% of cases. 2) For library usage, LLMs heavily favour top-ranked libraries such as NumPy, yet in up to 48% of cases, this usage is unnecessary and differs from the ground-truth solutions. 3) There is extremely low consistency between the programming languages and libraries LLMs recommend and those that they actually adopt, and the strong biases towards languages and libraries are universal across all the LLMs we study.
These findings reveal that widely used LLMs have notable diversity problems in their coding preferences, frequently straying from best practices and domain requirements. Moreover, such built-in biases may lead to homogenised code that limits creativity and niche solutions, and hinders the discoverability and growth of open-source software, which may ultimately weaken the effectiveness of LLM-based code generation.
The primary contributions of this paper are:
  • Presenting the first empirical study into LLM preferences when generating code, revealing a clear lack of diversity across LLMs when choosing programming languages and libraries, with significant preferences towards Python and established libraries.
  • Raising awareness of the need to measure and understand the diversity and preferences within LLM-generated code, providing a discussion of how this could impact the coding landscape and open-source software.
  • Releasing our code, datasets and full results publicly, to encourage further investigation in this area:
    https://github.com/itsluketwist/llms-love-python
The rest of the paper is organised as follows: Section 2 summarises work in related areas. Section 3 details the research questions, and the experiments conducted to answer them. Section 4 presents the results and answers to the research questions. Section 5 further discusses the findings and their broader impacts. Section 6 concludes the paper.
2. Related Work
To the best of our knowledge, this is the first paper to directly examine LLM preferences in programming languages and libraries for code generation. While no prior work specifically addresses this topic, several related problems have been explored in the literature.
Bias and Unfairness in LLMs: Whether an LLM has preferences for specific programming languages and libraries is related to bias and whether an LLM provides an unfair representation of the world. Bias within an LLMs natural language responses has been extensively studied (gallegosBiasFairnessLarge2024), but bias within any generated code is widely under explored. Most works focus on trustworthiness of the code (patelEvaluatingInContextLearning2024; wangInvestigatingDesigningTrust2024; mohsinCanWeTrust2024), or how LLMs can propagate harmful social biases within their code (liuUncoveringQuantifyingSocial2023; huangBiasTestingMitigation2023) - not whether the coding choices themselves include bias. A study has been conducted on the similarities between human and LLM preferences in task responses (liDissectingHumanLLM2024), but it does not go in depth on the preferences of code-generation.
LLM-based Code Generation: Code-generation in LLMs has been extensively studied (jiangSurveyLargeLanguage2024; zhangUnifyingPerspectivesNLP2023), yet the best way to evaluate model performance on code-generation tasks is an open problem (paulBenchmarksMetricsEvaluations2024a). Existing approaches typically use benchmarks that contain a dataset of programming problems and tests to check the functionality of the generated code (paulBenchmarksMetricsEvaluations2024a), although early benchmarks have shown to be limited in their scope (yadavPythonSagaRedefiningBenchmark2024a). Leveraging LLMs to perform there own code-evaluation has also been suggested (tongCodeJudgeEvaluatingCode2024), as well as the need to test LLMs in more realistic, repository-level, code-generation scenarios (jimenezSWEbenchCanLanguage2024a).
Library-oriented Code: Using external libraries to enhance one’s code is a key part of modern software development (somervilleSoftwareEngineeringGlobal2016). There are multiple studies that discuss the generation of library-oriented code (zanCERTContinualPretraining2022; liuCodeGen4LibsTwoStageApproach2023), including when private-libraries need to be used (zanPrivateLibraryOrientedCodeGeneration2023), and the API-recommendation problem (maCompositionalAPIRecommendation2024; wangSystematicEvaluationLarge2024) within LLMs. But none that discuss the impact that an LLMs preferences towards libraries may have on the coding landscape.
LLMs for Software Engineering: This paper argues that LLMs will impact how a developer makes decisions during software development. There are papers that discuss the problems of library selection (tanzilHowPeopleDecide2024; larios-vargasSelectingThirdpartyLibraries2020), and making implementation design choices (liangQualitativeStudyImplementation2023) by developers, without the use of LLMs. Using ChatGPT as a software librarian to aid these decisions has been suggested (latendresseChatGPTGoodSoftware2024), as well as how users’ distrust toward LLMs may impact the choices they make when using them as coding assistants (choudhuriWhatGuidesOur2024). An empirical study has also been done to analyse how an LLM can best benefit a software engineering project (rasnayakaEmpiricalStudyUsage2024).
3. Experimental Design
This section sets out the research questions (RQs), describing the methodology used and experiments designed to answer them.
3.1. Research Questions
  1. RQ1:
    Language Preference: What programming language preferences do LLMs exhibit in code generation? To answer this RQ, we analyse programming language preference in LLMs using widely-used benchmarks (Section 3.3.1) and more realistic project initialisation tasks (Section 3.3.2).
  2. RQ2:
    Library Preference: What library preferences do LLMs exhibit in code generation? To answer this RQ, we analyse coding library preference in LLMs using library-agnostic Python problems from BigCodeBench (Section 3.4.1) and more realistic project initialisation tasks (Section 3.4.2).
  3. RQ3:
    Internal Consistency: What is the internal consistency of LLMs between the programming languages and libraries they recommend for a task and those they actually use? To answer this RQ, we prompt LLMs to rank their recommended programming languages (Section 3.3.2) and Python libraries (Section 3.4.2). We then evaluate the alignment between their recommendations and actual choices using Kendall’s $\tau$ coefficient.
  4. RQ4:
    LLM Similarities: Do different LLMs have similar preferences when selecting programming languages and libraries? To answer this RQ, we examine the most frequently used languages and libraries across different LLMs in all experiments (Sections 3.3.13.3.23.4.13.4.2) and assess their similarity using Kendall’s $\tau$ coefficient to measure rank correlation across LLMs.
3.2. LLM Selection
A wide-range of LLMs are used for this study, to gain a broad understanding of the preferences of LLMs when generating code across models of different types. LLMs are chosen to vary in number of parameters, availability (open or closed source), and intended use-case (either general or code-specific). Eight LLMs were chosen, with full details given in Table 1.
We do not use Code Llama in this study. Although it is a popular LLM for code-generation, the markdown text it provides as output does not indicate the programming language being used, as described in Section 3.5.1. Therefore the automated extraction of languages from responses could not be performed reliably.
Model Usage: Each LLM is prompted with default settings and zero-shot prompting to reflect typical developer use. All prompts are run in fresh sessions to avoid bias from caching or leakage.
Table 1. Summary of LLMs used in this study.
Model Version Release Knowledge cut-off Size Open- source? Code model? GPT-4o (openaiGPT4TechnicalReport2024) gpt-4o-mini- 2024-07-18 July ’24 Oct. 2023 - GPT-3.5 (andrewpengGPT35TurboFinetuning) gpt-3.5-turbo- 0125 Nov. ’22 Sep. ’21 - Sonnet3.5 (anthropicClaude3Model2024) claude-3-5- sonnet-20241022 Oct. 2024 July 2024 - Haiku3.5 (anthropicClaude3Model2024) claude-3-5- haiku-20241022 Oct. ’24 July ’24 - Llama3.2 (grattafioriLlama3Herd2024) llama-3.2-3b- instruct-turbo Sep. ’24 Dec. ’23 3B Qwen2.5 (huiQwen25CoderTechnicalReport2024) qwen2.5-coder- 32b-instruct Nov. ’24 Mar. ’24 32B DeepSeekLLM (deepseek-aiDeepSeekLLMScaling2024) deepseek-llm- 67b-chat Nov. ’23 May ’23 67B Mistral7B (jiangMistral7B2023) mistral-7b- instruct-v0.3 May ’24 May ’24 7B
3.3. Experimental Setup for Studying Language Preferences
3.3.1. Benchmark Tasks
We begin by investigating the programming language preferences of LLMs when solving language-agnostic coding tasks from widely-used benchmarks.
Dataset Selection: We select datasets that are widely adopted to benchmark model performance and represent a range of different styles of problems. We require problem formulations to be described in natural language only, as any code in the prompt may bias the LLMs’ choices.
Additionally, to avoid bias, we select datasets that provide ground truth solutions in multiple programming languages, ensuring that LLMs are exposed to diverse language options rather than being influenced by a single dominant language. For some datasets, a few tasks contained language-specific references, these tasks were removed before sampling to further avoid bias. The selected six datasets can be categorised into three groups:
  1. (1)
    Basic: Short coding problems that an entry-level programmer could solve.
    • Multi-HumanEval (athiwaratkunMultilingualEvaluationCode2023): A multi-language version of the popular HumanEval (chenEvaluatingLargeLanguage2021) dataset originally published by OpenAI in 2021, to test code generation from the provided docstrings.
      Tasks: 967. Solution languages: C#, Go, Java, Javascript, Kotlin, Perl, PHP, Python, Ruby, Scala, Swift, Typescript.
    • MBXP (athiwaratkunMultilingualEvaluationCode2023): A multi-language version of the popular MBPP (austinProgramSynthesisLarge2021) dataset originally published by Google in 2021, containing simple problems that require short programs to be written as solutions.
      Tasks: 160. Solution languages: C#, Go, Java, Javascript, Kotlin, Perl, PHP, Python, Ruby, Scala, Swift, Typescript.
  2. (2)
    Real-world: Code generation tasks constructed from realistic coding situations that developers have faced. Because the problems are realistic, there are likely to be solutions available online in many languages.
    • AixBench (haoAixBenchCodeGeneration2022): A method-level code generation dataset built from comments found in public GitHub Java methods.
      Tasks: 161 total, 129 valid. Solution language: Java.
    • CoNaLa (yinLearningMineAligned2018): A code generation from natural language dataset built from coding problems found on StackOverflow.
      Tasks: 2879 total, 2255 valid. Solution language: Python.
  3. (3)
    Coding challenge: In-depth problems that typically require more thinking in order to solve. Sourced from online coding competition websites, they are language-agnostic by nature. Although the datasets provide solutions in specific languages, users are free to solve the problems in any language they choose, so there are solutions published online in many languages.
    • APPS (hendrycksMeasuringCodingChallenge2021): A large collection of problems from various open-access coding websites, categorised into 3 difficulty levels: introductory, interview, and competition.
      Tasks: 10000 total, 8918 valid. Solution language: Python.
    • CodeContests (doi:10.1126/science.abq1158): A competitive coding dataset initially used to train AlphaCode (CompetitiveProgrammingAlphaCode2024), labelled with the expected difficulty to solve.
      Tasks: 12664 total, 12638 valid. Solution languages: C++, Java, Python.
Prompt Strategy: We randomly select 200 problems from each dataset (spread proportionately across difficulty, for APPS and CodeContests) to use for the investigation. For each LLM, we generate three solutions per problem to reduce the effect of LLM randomness (sallouBreakingSilenceThreats2024), getting a fairer representation of their language preferences. Solutions are generated from LLMs using Prompt 1 (Table 2). An explanation is requested along with the solution because this helps to ensure that the response is in markdown format, and could therefore have the programming language automatically extracted (as described in Section 3.5.1).
Table 2. Prompts used when interacting with the LLMs.
1 Generate a code-based solution, with an explanation, for the following task or described function: <problem text> 2 Write the initial code for a <project description>. 3 List, in order, the best coding languages for the following task: Write the initial code for a <project description>. 4 Write a python function to solve: <problem text>. You should write self-contained python code. Choose, import and utilise at least one external library. 5 Write the initial code for a <project description>. You should write self-contained python code. Choose, import and utilise at least one external library. 6 List, in order, the best python libraries for the following task: Write the initial code for a <project description>.
3.3.2. Project Initialisation Tasks
To investigate language preferences in a more realistic setting, the LLMs are prompted to write the initial code for various new coding project descriptions. This is a typical use-case of LLMs by developers, as LLMs have been found to be particularly useful in the early stages of software-development projects (rasnayakaEmpiricalStudyUsage2024). Moreover, this stage is a natural decision point for selecting a programming language, where the LLM’s preferences may influence the user’s choice.
Task Selection: Our initial results imply that the LLMs have a general preference towards using Python. Therefore, for this task, project descriptions are chosen in areas where Python is not the best choice. Python is widely appreciated for its developer-friendly syntax and rapid prototyping capabilities, but it is not considered a high-performance language in all situations due to its interpretive nature and inefficient concurrency implementations (gavrilovaProsConsPython2023). For computationally intensive tasks, compiled languages like C, C++, and Rust are generally preferred because they provide greater memory control, lower execution overhead, and superior parallel processing capabilities (costanzoPerformanceVsProgramming2021).
To challenge LLM’s default language preference, we select the following five project descriptions that are representative of domains where high-performance, concurrency, or low-level control is crucial. These categories are commonly associated with languages optimised for speed and efficiency rather than ease of use:
  1. (1)
    Concurrency: "...high-performance web server to handle a large number of concurrent requests". High performance web servers require efficient thread management and concurrency control, which are areas where languages like Rust (with native async capabilities) or C++ (with multithreading support) outperform Python (bugdenSafetyPerformanceProminent2022).
  2. (2)
    Graphical interface: "...modern cross-platform
    application with a graphical user interface"
    . GUI applications benefit from native performance optimisations and low-latency rendering, making languages like C++ more suitable than Python (karczewskiPythonVsTechnology2021).
  3. (3)
    Low-latency: "...low-latency trading platform
    that will allow scaling in the future"
    . Financial trading systems demand minimal execution delays and precise memory management, areas where C++ and Rust are better than Python (bugdenSafetyPerformanceProminent2022).
  4. (4)
    Parallel processing: "...high-performance parallel task processing library". Effective parallelism requires fine-grained control over memory and execution threads, which compiled languages like C, C++, and Rust could provide (costanzoPerformanceVsProgramming2021).
  5. (5)
    System-level programming: "...command line
    application to perform system-level programming"
    . System programming typically relies on direct hardware interaction and efficient memory usage, making C, C++, and Rust better choices than Python (costanzoPerformanceVsProgramming2021).
Prompt Strategy: The project initialisation tasks are more open-ended, allowing for a wider range of possible solutions. To obtain a fairer representation of LLM preferences, we generate multiple iterations and analyse the distribution of languages used in the responses. For each LLM, 100 solutions are generated per problem using Prompt 2 (Table 2).
For RQ3: Internal Consistency, we also examine whether LLMs maintain consistency in their language recommendations — does the programming language they suggest for a task align with the one they actually use? To assess this, each model is prompted to provide an ordered list of recommended languages for each task using Prompt 3 (Table 2). This process is repeated three times to mitigate randomness, with an average of the rankings calculated to determine the final ranking per LLM.
3.4. Experimental Setup for Studying Library Preferences
To investigate library preferences, we focus on a single programming language to enable a more in-depth analysis. Python is chosen for library-related experiments due to its vast collection of open-source libraries [^1] The Python Package Index: https://pypi.org/, its easily parsable import syntax (as discussed in 3.5.1), and its prevalence in code-generation benchmarks.
3.4.1. Benchmark Tasks
We also wanted to investigate the Python library preferences of LLMs using widely-studied benchmarks that require the use of external libraries.
Dataset Selection: Whilst there are many library-based Python datasets, they are typically domain-specific, assessing how well an LLM uses certain libraries (such as DS-1000 (laiDS1000NaturalReliable2022)). We want to investigate model preferences when choosing coding libraries to solve library-agnostic tasks across a range of topics. BigCodeBench (zhuoBigCodeBenchBenchmarkingCode2024) contains 1140 Python coding problems across 7 domains (general, computation, visualisation, system, time, network and cryptography). Most problems require external libraries to be used when solving them, making it ideal for this use case.
BigCodeBench has been seeded from the ODEX dataset (wangExecutionBasedEvaluationOpenDomain2023a) (which originated from Stack Overflow) but has been perturbed and obfuscated to mitigate biases during evaluation. BigCodeBench was also released after the knowledge cut-off of the LLMs in this study, making it unlikely that these models have been exposed to its contents.
Dataset Processing: The problems in BigCodeBench have a natural language description of a function to be written, followed by the function output and the Python imports that should be used, therefore the problems needed to be updated to be library-agnostic. The function output and Python imports were removed from the end of the problem text, as they often had library-specific references. Then any remaining problems that still had library names in the text were removed. Full details of the dataset before and after processing are available in our GitHub repository [^2] Project repository: https://github.com/itsluketwist/llms-love-python.
Prompt Strategy: Each LLM solves BigCodeBench problems using Prompt 4 (Table 2), generating three solutions per problem to reduce randomness (sallouBreakingSilenceThreats2024) and ensuring a fair representation of the libraries used by each LLM.
3.4.2. Project Initialisation Tasks
For the same reasons discussed in Section 3.3.2, LLMs are asked to write the initial code for various library-agnostic project descriptions.
Task Selection: Project descriptions needed to be realistic for developers to want to build, require an external library to support its core functionality, and have many viable libraries that could be used. This will ensure that the responses make use of external libraries, and the LLMs have multiple valid options to choose from.
Project descriptions chosen are inspired by the groups of related libraries on Awesome Python [^3] Awesome Python: https://awesome-python.com/, ensuring there are multiple possible core libraries for each task. A selection of the available libraries are given for each task.
  1. (1)
    Database Object-Relational Mapping (ORM):
    "... database project with an object-relational
    mapping layer."
    Possible libraries: Django Models, peewee, Pony, pyDAL, SQLAlchemy.
  2. (2)
    Deep-learning: "... deep learning project
    implementing a neural network."
    Possible libraries: Caffe, Keras, PyTorch, scikit-learn, TensorFlow.
  3. (3)
    Distributed computing: "... distributed computing project." Possible libraries: Celery, Dask, Luigi, PySpark, Ray.
  4. (4)
    Web-scraper: "... web scraping and analysis
    library."
    Possible libraries: BeautifulSoup, lxml, MechanicalSoup, Scrapy, Selenium.
  5. (5)
    Web-server: "... backend API web server." Possible libraries: Django Rest, FastAPI, Flask, Pyramid, Starlette.
Prompt Strategy: The project initialisation tasks are again more open-ended. Therefore to ensure fairer representation of LLM preferences each task had 100 responses generated per model, using Prompt 5 (Table 2), to record the distribution of preferences.
For RQ3 Internal Consistency, the same process as in Section 3.3.2 was also followed to check the internal consistency of the LLMs when it comes to library recommendations. LLMs were asked for an ordered list of recommendations using Prompt 6 (Table 2), repeated three times to mitigate randomness, with an average of the rankings calculated to determine the final ranking per LLM.
3.5. Result Analysis
3.5.1. Response Data Extraction
In all experiments, LLM responses undergo an automatic process using regex pattern matching, to extract the programming language or libraries used. By default, models used in the study use markdown format when responding with code and natural language. In markdown, code blocks are denoted by a triple backtick followed by the programming language name (ExtendedSyntaxMarkdown2025), therefore a string search can be used to extract code blocks and programming languages.
For library experiments, if a Python code block has been extracted, a further string search is used to extract the imported libraries. There are two correct syntaxes for absolute imports in Python (nzomoAbsoluteVsRelative2025):
  1. (1)
    import <library name>
  2. (2)
    from <library name> import <object name>
3.5.2. Rank Correlation Analysis
We used Kendall’s $\tau$ coefficient to measure rank correlation for RQ3 Internal Consistency and RQ4 LLM Similarities. Kendall’s $\tau$ coefficient is a non-parametric measure of rank correlation that quantifies the ordinal association between two variables by comparing the number of concordant and discordant pairs in their rankings (kendallRankCorrelationMethods1990). We used the Kendall’s $\tau_{b}$ coefficient variant that allows for ties in both variables.
For RQ3 Internal Consistency, we use the results from the experiments investigating languages and libraries chosen for project initialisation tasks. We compare the LLMs’ rankings when asked for recommendations, with the ranking provided by the percentage of tasks that the language or library is used in.
For RQ4 LLM Similarities, we used the results from all of the experiments. We compare the languages or libraries used by each LLM, against those used by each other LLM, to measure similarity in each scenario.
Table 3. Language Preferences, Benchmark Tasks (RQ1). Programming languages used by LLMs when solving problems from widely-used benchmarks. Column “Dataset” shows the name of the benchmark, as well as the programming languages for which the ground truth solutions are provided in each benchmark. For each LLM and dataset, the preferred languages (l) are given, along with the percentage (p) of dataset problems that have a solution in that language, and the total count of different languages used (when necessary). The most-used language in each case is in bold.
Dataset GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b l p l p l p l p l p l p l p l p Multi-HumanEval (10 languages*) Python - 100.0% - Python - 100.0% - Python JavaScript 98.8% 1.2% Python Total used 99.4% 4 Python - 100.0% - Python - 100.0% - Python - 100.0% - Python Total used 100.0% 3 MBXP (10 languages*) Python - - - 100.0% - - - Python JavaScript - - 100.0% 1.0% - - Python JavaScript Java C 97.0% 7.5% 1.0% 0.5% Python Java C++ JavaScript 99.5% 2.5% 2.0% 2.0% Python - - - 100.0% - - - Python - - - 100.0% - - - Python - - - 98.5% - - - Python JavaScript - - 99.0% 0.5% - - AixBench (Java) Python Java JavaScript Total used 75.2% 27.1% 10.9% 7 Python Java JavaScript Total used 64.3% 38.8% 10.1% 7 Java Python JavaScript Total used 53.5% 38.8% 18.6% 9 Java Python JavaScript Total used 59.7% 55.0% 20.2% 12 Python Java C# Total used 75.2% 18.6% 3.1% 8 Python Java JavaScript Total used 68.2% 26.4% 3.1% 5 Python Java JavaScript C# 78.3% 18.6% 2.3% 0.8% Python Java JavaScript Total used 62.0% 24.8% 14.0% 9 CoNaLa (Python) Python JavaScript Java Total used 99.0% 6.0% 4.5% 7 Python JavaScript Java C++ 98.5% 2.5% 1.5% 0.5% Python JavaScript Java Total used 98.0% 12.5% 3.5% 9 Python JavaScript Java Total used 99.0% 15.5% 13.0% 11 Python JavaScript - - 99.0% 1.0% - - Python JavaScript Java Total used 98.0% 3.0% 2.0% 5 Python JavaScript Java C++ 99.0% 1.5% 1.0% 0.5% Python JavaScript Java Total used 97.5% 4.0% 2.5% 5 APPS (Python) Python JavaScript Ruby R 99.5% 1.0% 0.5% 0.5% Python JavaScript Ruby Java 99.5% 1.5% 0.5% 0.5% Python C++ JavaScript Total used 93.5% 10.0% 5.5% 5 Python C++ JavaScript Total used 93.5% 7.5% 4.5% 6 Python Ruby Java C 98.0% 0.5% 0.5% 0.5% Python C++ Ruby R 98.0% 1.5% 0.5% 0.5% Python Ruby - - 98.5% 0.5% - - Python Ruby Java R 98.5% 0.5% 0.5% 0.5% CodeContests (C++, Java, Python) Python - - 100.0% - - Python C++ - 98.5% 2.0% - Python C++ Java 96.5% 19.5% 0.5% Python C++ - 94.0% 15.0% - Python C++ JavaScript 97.5% 2.0% 0.5% Python C++ - 97.0% 6.0% - Python - - 96.5% - - Python C++ JavaScript 97.5% 2.0% 0.5% All datasets Python Java JavaScript C# Total used 96.8% 4.0% 2.6% 0.5% 11 Python Java JavaScript C++ Total used 95.1% 5.0% 2.1% 0.6% 8 Python Java JavaScript C++ Total used 89.8% 7.3% 7.1% 6.2% 13 Python Java C++ JavaScript Total used 92.0% 10.3% 7.8% 6.6% 14 Python Java C++ JavaScript Total used 96.1% 2.3% 0.6% 0.5% 9 Python Java C++ JavaScript Total used 94.9% 3.5% 1.7% 0.9% 8 Python Java JavaScript C# Total used 96.1% 2.4% 0.6% 0.1% 6 Python Java JavaScript C# Total used 94.1% 3.5% 2.6% 0.6% 13
  • * Multi-HumanEval and MBXP benchmarks contain solutions in the following languages: C#, Go, Java, JavaScript, Kotlin, Perl, PHP, Python, Ruby, Scala, Swift, TypeScript.
4. Results
This section introduces the results. Note that it is expected that the percentages in the results do not add up to 100% because LLMs were prompted multiple times per task and sometimes responded with different languages or libraries. Additionally, the LLMs would sometimes provide a response without code; or would provide a response that contained code in multiple languages or that imported multiple libraries.
4.1. RQ1: Language Preference
This RQ aims to assess LLMs’ preferences when choosing programming languages for code-generation tasks.
4.1.1. Preferences for benchmark tasks
Table 3 shows the languages used by LLMs when solving problems from widely-used benchmarks, 200 problems were used per benchmark and each problem had three solutions generated. We observe a clear preference towards using Python for all LLMs and in all datasets.
For each dataset apart from AixBench, all LLMs gave solutions in Python for at least 93.5% of problems, the highest percentage of problems solved by a non-Python language was considerably lower at 19.5%. AixBench had a much more even distribution, with Python solutions given for only 53.5-78.3% of problems, with the second most-used language, Java, having solutions for 18.6-59.7% of problems. AixBench problems solved by the Claude models (Haiku3.5 and Sonnet3.5) are the only instances where Python was not the most-used language.
The “Real-world” datasets (AixBench and CoNaLa) had up to 12 different languages used in their solutions, significantly more than the other datasets - which had at most 5 different languages used, showing that models opted for more suitable languages in real-world situations. In general the range of programming languages that LLMs choose to use is limited, hundreds of programming languages get contributions on GitHub every year (TopProgrammingLanguages2022), but LLMs only choose to use 6-14 different ones.
We suspect this heavy preference for Python stems from LLMs being created with a focus on achieving state-of-the-art results on widely-used benchmarks, the vast majority of which are Python based, causing their training data to be saturated with Python code. When investigating what datasets to use for this study, it was clear that there was a lack of diversity in the datasets available and it was difficult to find suitable options with no link to Python. Whilst all datasets used have language-agnostic problem formations, some have been seeded from Python (Multi-HumanEval, MBXP, CoNaLa), or have Python solutions published (APPS, CodeContests). AixBench is the only dataset with no link to Python, which is likely why it is the closest to favouring another language.
Table 4. Language Preferences, Project Initialisation Tasks (RQ1). Programming languages chosen by LLMs for initial project code in scenarios where Python is not the most suitable option. For each LLM and project, each language (l) used by the LLM is given, along with the percentage (p) of responses that used the language and the rank (r) assigned to that language from the LLM. The most-used language in each case is in bold.
Language Task GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b l p r l p r l p r l p r l p r l p r l p r l p r Concurrency JavaScript Python Go 73% 28% 2% #3 #8 #1 JavaScript Python - 99% 1% - #5 #6 - JavaScript - - 100% - - #5 - - Python - - 100% - - - - - Go - - 100% - - #2 - - Python - - 100% - - #12 - - JavaScript Python - 94% 6% - #5 #4 - JavaScript - - 100% - - - - - Graphical interface JavaScript Dart Python C++ 72% 38% 14% 3% #1 #5 #3 #12 JavaScript C++ Python - 64% 29% 11% - #1 - #2 - Python Dart - - 99% 2% - - #3 #1 - - Python JavaScript TypeScript - 81% 19% 2% - #3 #2 - - Python - - - 100% - - - #5 - - - JavaScript Dart - - 100% 100% - - - #6 - - Python - - - 100% - - - #2 - - - Dart JavaScript - - 81% 48% - - #2 #1 - - Low-latency Python - - - 100% - - - #5 - - - Python Java JavaScript C++ 96% 2% 2% 1% #3 #2 - #1 Java C++ Python - 63% 30% 7% - #2 #1 - - Python - - - 100% - - - #5 - - - Python - - - 100% - - - #5 - - - Python - - - 100% - - - #5 - - - Python - - - 100% - - - #3 - - - Python none - - 53% 47% - - #3 - - - Parallel processing Python C++ - 99% 1% - #5 #1 - Python C++ Java 80% 11% 9% #5 #1 #4 Python C++ - 96% 4% - - #2 - Python - - 100% - - - - - Python - - 100% - - #6 - - C++ - - 100% - - #1 - - Python - - 100% - - #3 - - Python - - 100% - - #6 - - System-level programming Python C 81% 23% #5 #1 C Python 63% 37% #1 #5 C Python 50% 50% #1 #6 Python - 100% - - - Python - 100% - #6 - C - 100% - #1 - C++ - 100% - #2 - C++ C 51% 49% #1 #1
4.1.2. Preferences for project initialisation tasks
Table 4 shows the languages used when the LLMs wrote the initial code for different project descriptions over 100 iterations. The LLMs should favour highly-performant languages—such as C, C++ and Rust (costanzoPerformanceVsProgramming2021)—for the chosen tasks.
Even for project initialisation tasks where Python is typically not the best choice, there is still a clear preference from LLMs towards using it. Across models and tasks, Python is the preferred language in 23/40 cases even though it was given a low recommendation each time. Haiku3.5, Llama3.2 and DeepSeek were the worst offenders here, heavily using Python for 4/5 of the tasks. Javascript is the next most-used but still only the preferred choice in 8/40 instances. This result is consistent with the preference towards Python when solving benchmark tasks. We suspect that the trend towards favouring Python and Javascript comes from models being built to prefer more user-friendly languages.
LLama3.2 and Qwen2.5 are the models with the strongest preferences, using the same languages in every response. Haiku3.5 and DeepSeek only show slightly more diversity, having variability in the languages chosen for a single project description. The GPT models had the most variability in their responses, using a range of languages across their responses to each task - but they still only used at most four different languages for a single project.
Answer to RQ1: All LLMs exhibit a significant bias toward Python as their preferred programming language. Specifically, for benchmark tasks, LLMs use Python for 90–97% of all problems. When generating initial project code where Python is not a suitable language, Python remains the dominant choice, appearing as the most-used language in 58% of cases.
Figure 1. Library Preferences, Benchmark Tasks (RQ2). Libraries used by LLMs when solving problems from the BigCodeBench dataset. For each LLM, the most-used libraries are given with the percentage of problems they were imported for, along with total unique libraries used. All libraries not shown are imported for less than 2% of problems.
[…]
A bar chart showing library preferences of LLMs for benchmark tasks, detailing the most-used libraries and their usage percentages.
4.2. RQ2: Library Preference
This RQ aims to assess the LLMs’ preferences when choosing libraries for code-generation tasks.
4.2.1. Preferences for benchmark tasks
Figure 1 shows the libraries used by models when solving 526 library-agnostic problems from BigCodeBench, with three solutions generated per problem. All LLMs show similar preferences across the problems, with a similar distribution in the most-used libraries.
In each case, the top 3 libraries are identical (NumPy, pandas and Matplotlib) and heavily preferred, with a gap of at least 22% between themselves and the fourth most-used library. It is unsurprising that these libraries are the most-preferred due to a large number of the contained problems being in the computation and visualisation domains, but they are often imported when they are not necessary. For example, the most used library, NumPy, is imported in solutions to 192 of the 301 problems where the ground truth solution does not contain it at all - showing a strong preference to use it over alternative options for any task.
The range of Python libraries that are available for download is expansive - in the month of January 2025, over 7000 different open-source libraries had over 100,000 downloads (TopPyPIPackages2025) (and this is still just a fraction of those available). However, looking at the full results from the figure, all models use similarly low numbers of unique libraries, ranging from 32 (Llama3.2) to 39 (Haiku3.5). Whilst not all of those libraries are suitable for the given problems, using less than 40 over a varied range of problems shows a severe lack of diversity.
Table 5. Library Preferences, Project Initialisation Tasks (RQ2, RQ3). Libraries used by LLMs when writing initial project code. For each LLM and project, the preferred libraries (l) are given, along with the percentage (p) of responses that used the library and the rank (r) assigned to that library from the LLM. Libraries considered to provide core functionality are marked with a * and are listed first.
Library Task GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b l p r l p r l p r l p r l p r l p r l p r l p r Database ORM SQLAlchemy* models database - 100% 11% 3% - #1 - - - SQLAlchemy* - - - 100% - - - #1 - - - SQLAlchemy* - - - 100% - - - #1 - - - SQLAlchemy* Pydantic dotenv Total used 100% 3% 3% 5 #1 - - - SQLAlchemy* models db.py - 100% 96% 1% - #1 - - - SQLAlchemy* TensorFlow - - 99% 1% - - #1 - - - SQLAlchemy* - - - 100% - - - #1 - - - SQLAlchemy* - - - 100% - - - #1 - - - Deep- learning TensorFlow* scikit-learn* PyTorch* TorchVision* NumPy Matplotlib 97% 4% 3% 2% 76% 64% #1 #10 #2 - - - TensorFlow* PyTorch* Keras* scikit-learn* TorchVision* NumPy 85% 7% 6% 2% 1% 47% #1 #2 #3 #4 - #5 PyTorch* TorchVision* scikit-learn* Matplotlib NumPy - 100% 93% 5% 96% 11% - #2 - #5 #8 #4 - scikit-learn* TensorFlow* PyTorch* TorchVision* NumPy Matplotlib 82% 75% 25% 14% 86% 26% - #1 #2 - - - TensorFlow* scikit-learn* NumPy Matplotlib - - 100% 91% 100% 91% - - #3 #7 #1 #8 - - TensorFlow* - - - - - 100% - - - - - #1 - - - - - TensorFlow* - - - - - 100% - - - - - #1 - - - - - Keras* scikit-learn* NumPy - - - 100% 44% 100% - - - #3 #4 - - - - Distributed computing Dask* MPI4py* Ray* NumPy Total used 73% 5% 2% 28% 8 #1 #8 #2 - - MPI4py* Dask* NumPy Joblib - 33% 31% 6% 1% - #5 #1 - - - Dask* Ray* Celery* Redis Total used 24% 20% 1% 49% 8 #2 #1 #4 - - Dask* Ray* Celery* NumPy Total used 89% 4% 2% 71% 6 #1 #2 #4 - - Dask* NumPy - - - 100% 55% - - - #1 #9 - - - Ray* - - - - 100% - - - - #2 - - - - - - - - - - - - - - - - - - - Dask* NumPy - - - 100% 100% - - - #1 - - - - Web-scraper BS4* Requests* pandas - 100% 100% 27% - #2 #1 #5 - BS4* Requests* pandas - 100% 100% 13% - #1 #4 #3 - BS4* Requests* pandas - 100% 100% 100% - #2 #1 #5 - BS4* Requests* pandas Total used 100% 100% 100% 6 #8 #1 #6 - BS4* Requests* pandas - 100% 100% 100% - #1 #3 #8 - BS4* Requests* pandas - 100% 100% 100% - #2 #1 #5 - BS4* Requests* - - 100% 100% - - #1 #2 - - BS4* Requests* pandas - 100% 100% 100% - #2 #1 #3 - Web-server Flask* FastAPI* Flask-REST Flask-Cors Total used 98% 2% 19% 15% 7 #1 #2 - - - Flask* - - - - 100% - - - - #1 - - - - FastAPI* Flask* Pydantic Uvicorn Flask-Cors 48% 39% 48% 48% 14% #1 #2 #12 #14 - Flask* FastAPI* Flask-Cors Flask-SQLA Total used 69% 23% 52% 31% 12 #2 #1 - - - Flask* Flask-SQLA - - - 100% 100% - - - #1 - - - - Flask* - - - - 100% - - - - #1 - - - - Flask* - - - - 100% - - - - #1 - - - - Flask* - - - - 100% - - - - #1 - - - -
  • Note some library names have been shortened in the table: Flask-REST is Flask-RESTful, Flask-SQLA is Flask-SQLAlchemy, BS4 is BeautifulSoup.
Case Analysis: It is clear from the results that all LLMs show a preference for established libraries. To understand the extent of this bias, we conduct a case analysis on libraries used in different domains, to see if there are any high-quality libraries that are neglected by LLMs. In particular, we use GitHub star amount and growth over time as simple objective metrics for library quality, and look for alternative libraries that are showing signs of higher quality but have lower usage rates.
The results are shown in Table 6, giving GitHub statistics and usage rate (the percentage of problems where a solution included that library) for studied libraries, where we observe that high-quality alternatives exist in the computation and visualisation domains.
For computation problems, pandas is heavily favoured. Polars is a newer library with similar functionality (including usage of DataFrame objects) and twice as fast GitHub star growth, but is not used at all. Similarly, Matplotlib and seaborn are heavily used for visualisation problems. Plotly is again newer, has additional functionality and signs of higher potential, but is used for only one problem and by three LLMs.
Table 6. Case Analysis. Statistics about libraries studied in case analysis, and their repositories (data queried from GitHub on 5th December 2024). For “Domain” rows, the usage rate gives the percentage of problems where that library was imported; for “Project” rows, the usage rate gives the percentage of solutions where that library was imported. Growth is given as GitHub stars per year, largest growth and usage rate is in bold.
Library GitHub Repository Repository Creation Stars Growth Usage Rate $\downarrow$ Computation pandas pandas-dev/pandas Aug. 2010 44.0k 3.1k 58% polars pola-rs/polars May 2020 30.7k 6.7k 0% Visualisation Matplotlib matplotlib/matplotlib Feb. 2011 20.4k 1.5k 57% seaborn mwaskom/seaborn June 2012 12.6k 1.0k 17%  Domain Plotly plotly/plotly.py Nov. 2013 16.4k 1.5k 0.4% Distributed computing Dask dask/dask Jan. 2015 12.7k 1.3k 52% Ray ray-project/ray Oct. 2016 34.3k 4.2k 16% Celery celery/celery Apr. 2009 25.0k 1.6k 0.4% Web-server Flask pallets/flask Apr. 2010 68.2k 4.7k 88%  Project FastAPI fastapi/fastapi Dec. 2018 30.7k 13.1k 9%
4.2.2. Preferences for project initialisation tasks
Table 5 shows the libraries imported when the LLMs wrote the initial code for different Python project descriptions over 100 iterations.
We observe that DeepSeek and Mistral7b have the least strong preferences, using the same libraries in every response. Llama3.2 and Qwen2.5 are only slightly better, using the same core libraries in every response and only varying the supporting libraries. The closed-source GPT and Claude LLMs showed much more varied preferences, with differing core libraries in 3/5 tasks: “Deep-learning”, “Distributed computing” and “Web-server”.
Notably, there are multiple instances where a library was imported for a project that would not require it: Qwen2.5 imported TensorFlow for the “Database ORM” task; and multiple models imported NumPy and pandas for the “Distributed computing” task. This over-usage of various data-science libraries is consistent with the results for benchmark tasks and reinforces the bias towards their usage. These preferences could be correlated to the recent growth in Python usage that has been attributed to the growth in AI and data-science based code (staffOctoverseAILeads2024a).
Case Analysis: The results once again show a preference for more established libraries. We conduct further case analysis (as described in Section 5), to show how LLMs rarely use high-quality alternatives to their preferred libraries for project initialisation tasks. The results are shown in Table 6, giving GitHub statistics and usage rate (the percentage of responses using that library) for studied libraries, where we observe that high-quality alternatives exist for the “Web-server” and “Distributed computing” tasks.
For the “Web-server” task the potential Python frameworks are plentiful, Flask is a traditional option and was widely used by LLMs. FastAPI is much newer and has grown its GitHub stars almost three times as quickly - yet only three LLMs used it. For the “Distributed computing” task, Dask is the library favoured by most LLMs. Ray and Celery both have similar functionality along with more stars and faster growth, but they were used minimally, and by only four LLMs.
Answer to RQ2: There is a significant lack of diversity in the coding libraries chosen by LLMs. LLMs heavily favour data-science libraries, even when they are not necessary. For 48% of the problems where NumPy was imported, its usage is unnecessary and differs from the ground-truth solutions. LLMs also strongly favor older, well-established libraries, often overlooking newer, highly-rated alternatives, as shown in our four case studies, where usage rate differences range from 36% to 79%.
4.3. RQ3: Internal Consistency
This RQ aims to assess whether the programming languages or libraries LLMs recommend for a task are consistent with what they actually use. To answer this RQ, we ask the LLMs to rank the languages or libraries they would recommend for each project initialisation task, and used Kendall’s $\tau$ coefficient to compare this ranking to what they actually use. Results are shown in Table 7.
Table 7. Internal Consistency, Project Initialisation Tasks (RQ3). Kendall’s $\tau$ correlation between LLMs’ recommendations and their actual choices (along with the corresponding $p$-value). Values near 1.0/-1.0 indicate strong agreement/disagreement. Correlation values that are statistically significant ($p$-values $<$ 0.05) are in bold.
GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b Concurrency $\tau$ 0.31 -0.33 -0.35 -0.51 0.36 -0.41 0.04 -0.49 $p$ 0.20 0.34 0.38 0.12 0.32 0.11 0.88 0.12 Graphical interface $\tau$ 0.28 0.27 0.06 0.06 0.05 -0.55 0.36 0.60 $p$ 0.25 0.39 0.87 0.87 0.86 0.10 0.28 0.17 Low-latency $\tau$ 0.09 0.16 0.30 -0.35 -0.35 -0.35 0.00 -0.20 $p$ 0.75 0.63 0.42 0.38 0.38 0.38 1.00 0.57 Parallel processing $\tau$ 0.35 -0.32 -0.20 -0.53 -0.58 0.53 0.25 -0.36 $p$ 0.17 0.45 0.57 0.13 0.14 0.13 0.38 0.32  Language Tasks System-level programming $\tau$ 0.34 0.26 0.00 -0.58 -0.36 0.50 0.35 0.71 $p$ 0.20 0.50 1.00 0.14 0.32 0.13 0.22 0.07 Database ORM $\tau$ 0.45 0.58 0.63 0.63 0.47 0.53 0.45 0.63 $p$ 0.12 0.14 0.16 0.16 0.12 0.13 0.12 0.16 Deep-learning $\tau$ 0.14 0.57 -0.06 0.02 0.28 0.41 0.41 0.30 $p$ 0.55 0.02 0.83 0.93 0.30 0.11 0.11 0.31 Distributed computing $\tau$ 0.41 0.35 0.60 0.60 0.19 0.33 - 0.44 $p$ 0.02 0.17 0.02 0.01 0.45 0.19 - 0.11 Web-scraper $\tau$ 0.49 0.47 0.42 0.05 0.36 0.55 0.60 0.73 $p$ 0.01 0.17 0.01 0.82 0.17 0.04 0.04 0.03  Library Tasks Web-server $\tau$ 0.31 0.50 0.14 0.02 0.50 0.53 0.45 0.58 $p$ 0.17 0.13 0.50 0.93 0.13 0.13 0.12 0.14
From the table, we observe that there is extremely low internal consistency between the programming languages or libraries recommended by LLMs and LLMs’ final choices. In particular, for language preferences, none of the correlations are statistically significant. For library preferences, there are only a few significant correlations for tasks “Distributed computing” and “Web-scraper”, probably due to the fact that there are more widely accepted and commonly used libraries for these two categories.
The reason for such inconsistency is that the LLMs we study do not deliberately plan their responses in the way a human programmer would. Instead, they generate text token by token, predicting the most likely next word based on patterns in their training data. This means that when actually generating code, LLMs may default to the most frequently seen programming language or libraries in their training data (jonesCapturingFailuresLarge2022).
The inconsistency we observe highlights potential issues in the decision-making process of LLMs, raising concerns about their reliability in guiding developers toward optimal language and library choices. In Section 5.2.2, we discuss whether chain-of-thought (CoT) or other CoT-based LLMs such as DeepSeek-R1 mitigate this issue.
Answer to RQ3: There is low consistency between the programming languages and libraries LLMs recommend and those that they actually adopt, with the majority of correlations not statistically significant. For project initialisation tasks, LLMs contradict their own language recommendations 83% of the time.
4.4. RQ4: LLM Similarities
This RQ looks at whether preferences for languages and libraries are similar across different LLMs. To answer this RQ, Kendall’s $\tau$ coefficient is calculated for each previous experiment, correlating the preferences from each pair of LLMs. Table 8 shows the results for benchmark tasks.
From the table, we can observe that there is a median coefficient of 0.54 (range 0.40-0.67) across all LLM pairs for languages used across benchmark tasks, with only three results not being statistically significant. For libraries used on BigCodeBench tasks, correlation results for all LLM pairs are statistically significant, with a median coefficient of 0.53 (range 0.40-0.65). This indicates that all LLMs have similar preferences when solving benchmark tasks.
For project initialisation tasks, the correlation between preferences is much less clear. Only 16% of coefficients between LLMs have statistical significance ($p$-value $>$ 0.05), and 13% are undefined due to an exact match of a single choice of technology.
We assume the similarity across models reflects shared training data sources, like public GitHub repositories, while differences arise from data selection and training variations. These differences appear more pronounced for the open-ended project initialisation tasks, hence the lack of significance in the coefficients.
Answer to RQ4: Different LLMs have a notable alignment on their preferences for programming languages and libraries when solving benchmark tasks, with an overall median $\tau$ coefficient of 0.53 (range 0.40-0.67). There is no clear evidence of alignment on language and library preferences for project initialisation tasks - only 16% of $\tau$ coefficients have statistical significance.
Table 8. LLM Similarities, Benchmark Tasks (RQ4). Kendall’s $\tau$ correlation between different LLMs’, comparing the languages and libraries they use for benchmark tasks. All correlations given are statistically significant ($p$-values $<$ 0.05). Values near 1.0/-1.0 indicate strong agreement/disagreement.
GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b GPT-4o - 0.66 0.58 0.49 0.44 - 0.45 0.52 Language Tasks GPT-3.5 0.58 - 0.58 0.55 0.57 - 0.62 0.67 Sonnet3.5 0.49 0.62 - 0.50 0.60 - 0.47 0.51 Haiku3.5 0.50 0.65 0.60 - 0.56 0.40 0.57 0.43 Llama3.2 0.65 0.57 0.50 0.55 - 0.51 0.53 0.65 Qwen2.5 0.57 0.62 0.54 0.61 0.51 - - 0.44 DeepSeek 0.51 0.62 0.44 0.52 0.55 0.47 - 0.65 Mistral7b 0.44 0.55 0.40 0.44 0.49 0.43 0.50 - Library Tasks
5. Discussion
This section further discusses our findings and provides extended analysis on techniques that may impact LLM preferences.
5.1. Advantages & Disadvantages of LLMs’ Code Preferences
Our results show that LLMs exhibit strong preferences — and thus a lack of diversity — in their choice of programming languages and libraries. While we have previously discussed the disadvantages of these preferences when LLMs select inappropriate options, here we consider both the advantages and disadvantages of such preferences in cases where multiple choices, including the ones favoured by LLMs, are equally acceptable.
Advantages: LLMs often favor widely adopted technologies, which are typically mature, well-documented, and supported by large communities. This can enhance the user experience by making the generated code easier to understand, extend, and integrate. The use of established languages and libraries also increases the likelihood of generating robust and secure code, as these technologies have usually been thoroughly tested in real-world applications.
Disadvantages: However, these strong biases can lead to code homogeneity, limiting creative solutions and preventing LLMs from suggesting more specialised or optimal tools for specific tasks. This can be already seen multiple times in this study, most notably when Rust was not used a single time by any models for 5 different high-performance tasks (Table 4).
Bias toward mainstream technologies can also result in unnecessary or suboptimal dependencies, potentially causing performance issues. This is evidenced by the willingness of all models in this study to import and use data-science libraries across a range of tasks where they are not strictly necessary (Table 5).
Additionally, the lack of diversity in the preferences of LLMs will also likely lead to inadequate discoverability for open-source software. In Section 4.2 we present four use-cases where LLMs show a heavy preference towards older libraries over newer high-quality alternatives. If these biases persist, they risk reinforcing the dominance of a limited set of tools, stifling competition and innovation within the open-source ecosystem.
In-particular, this will severely limit the motivation to develop new open-source software. Open-source software is indispensable - evidenced by the reliance on Kubernetes and Docker for modern software development (NeedDockersKubernetes2024) and how 60% of websites are hosted using open-source web servers (UsageStatisticsMarket2024). Therefore, it is crucial for LLMs to contribute to, rather than hinder, open-source software growth.
5.2. Extended Analysis
This section introduces our investigations into how varying temperature or using CoT prompting may affect LLM preferences.
5.2.1. Varying Temperature
Temperature is a model parameter that affects the variability and randomness of responses, sometimes considered the creativity parameter (peeperkornTemperatureCreativityParameter2024). We conduct an initial investigation into how adjusting temperature may allow LLMs to diversify their choice of programming language during project initialisation tasks.
We repeat the experiment described in Section 3.3.2 for a single LLM (gpt-4o-mini-2024-07-18), whilst varying the temperature. The OpenAI API accepts temperatures from 0.0 to 2.0, but larger temperatures can lead to unreliable parsing of responses (renzeEffectSamplingTemperature2024), so the following temperature values were used: 0.0, 0.5, 1.0 and 1.5.
The results are shown in Table 9. Altering temperature has a clear but minimal impact - in each instance, a higher temperature led to the most-used language being used in a lower percentage of responses, along with a wider variety of languages. On average, the usage frequency of the most-used language will drop by 13.46% when increasing the temperature from 1.0 to 1.5. Interestingly, a temperature of 0.0 does not guarantee the most-used language to remain consistent.
Table 9. Varying Temperature. Languages used by GPT-4o for project initialisation tasks, when varying temperature. For each temperature (t) and project, the preferred languages (l) are given, along with the percentage (p) of responses that used that language, and a count of the total used (if necessary). The most-used language in each case is in bold.
Language Task t = 0.0 t = 0.5 t = 1.0 t = 1.5 l p l p l p l p Concurrency JavaScript Python Go 73% 27% 1% JavaScript Python Go 63% 36% 1% JavaScript Python Go 68% 32% 3% JavaScript Python Total used 58% 39% 6 Graphical interface JavaScript Dart Total used 84% 41% 4 JavaScript Dart Total used 94% 50% 4 JavaScript Dart Total used 76% 48% 5 JavaScript Dart Total used 62% 30% 6 Low-latency Python C++ 100% 1% Python - 100% - Python - 100% - Python JavaScript 65% 1% Parallel processing Python C++ 99% 1% Python - 100% - Python C++ 99% 1% Python C++ 97% 4% System-level programming Python C - 87% 16% - Python C - 93% 9% - Python C - 89% 17% - Python C Go 83% 17% 1%
5.2.2. Chain-of-thought Prompting
Chain-of-thought (CoT) prompting has been shown to elicit reasoning in LLMs (weiChainofthoughtPromptingElicits2022), and may allow them to have better internal consistency for the experiments in this study, recommending more suitable and higher-quality languages and libraries instead of what is most prevalent in their training data. CoT reasoning can also be used to train LLMs to reason–such as for recent LLMs like DeepSeek-R1 and OpenAI o1–or when prompting the LLMs for a response by including the phrase “think step by step” (kojimaLargeLanguageModels2022).
We conduct an initial investigation into how CoT may help LLMs with language choice. We repeat the experiment described in Section 3.3.2 for a single LLM (gpt-4o-mini-2024-07-18), asking for project initialisation code. We append the prompts with ”Think step by step about which programming language you should use and why.”, to elicit CoT-like reasoning from the LLM (kojimaLargeLanguageModels2022).
Table 10. Chain-of-thought Prompting. Languages used for project initialisation tasks by GPT-4o when using CoT prompting. The languages used are given, with the percentage of responses that used the language, and the rank assigned to the language by the LLM. The top-ranked language in each case is in bold.
Language Task Language Usage $\downarrow$ Rank
Concurrency Go JavaScript 91% 10% #1 #3
Graphical interface JavaScript Dart Python 69% 16% 16% #1 #5 #3
Low-latency C++ Go Python Rust Java JavaScript 89% 7% 5% 2% 1% 1% #1 #4 #5 #3 #2 #9
Parallel processing Rust Go C++ Python 61% 20% 18% 1% #2 #3 #1 #5
System-level programming C Rust Python 88% $1$3 #5
The results are shown in Table 10. Internal consistency has improved but is still not perfect, with the top-ranked programming language now being the most used in 4/5 instances. The biggest difference can be seen in the “Low-latency” and “Parallel processing” tasks, where usage of Python has decreased by 95% and 98% respectively.
6. Conclusion
This paper provides the first empirical investigation into LLM preferences when generating code, revealing a significant lack of diversity in programming language and library choices. Our findings demonstrate that LLMs exhibit a strong bias towards Python, preferring it for both benchmark tasks and project initialisation tasks, even when there are more suitable choices. Models also show a lack of internal consistency regarding programming language recommendation, frequently suggesting alternative languages while defaulting to Python when actually writing the code.
Similarly, LLMs display a limited range when choosing coding libraries, favouring well-established options over high-quality alternatives. This lack of diversity raises concerns about the homogenisation of the coding landscape due to LLMs becoming a significant contributor, and the potential negative impact on open-source discoverability.
These results highlight the need for further research in this area, ensuring that future LLMs contribute positively to the software ecosystem rather than reinforcing existing biases. We hope to inspire future work on improving the diversity of LLM-generated code–via techniques such as prompt engineering, adaptive decoding, reinforcement learning, and retrieval-augmented generation–as well as investigating other potential biases in LLM-generated code, inlcuding preferences for coding paradigms, architectures, data structures, and typing methods.
[…]